OpenStudy (anonymous):

suppose that u and v are linearly independent vectors. show that 2u+3v and u+v are linearly independent

OpenStudy (zarkon):

suppose that they are dependent and derive a contradiction

OpenStudy (anonymous):

What do you mean?

OpenStudy (zarkon):

assume there exists numbers \(a\) and \(b\) such that \[a(2u+3v)+b(u+v)=0\] then show that this leads so a contradiction with the fact that u and v are independent

OpenStudy (zarkon):

with a,b not both zero

OpenStudy (anonymous):

what do you mean by contradiction though? I have zero idea :\

OpenStudy (zarkon):

get a result that is not possible.. [for example 0=1 is a contradiction]

OpenStudy (zarkon):

take \[a(2u+3v)+b(u+v)=0\] show that a and b have to be zero

OpenStudy (anonymous):

So like I just sub in random numbers to see?

OpenStudy (zarkon):

no... assume \(a,b\neq\)0 (or at least on of them is non-zero) \[a(2u+3v)+b(u+v)=0\] \[\Rightarrow (2a+b)u+(3a+b)v=0\] since \(u,v\) are independent then \(2a+b=0\) and \(3a+b=0\) now show that a,b are both zero

OpenStudy (anonymous):

well b=-3a and -2a so wouldn't that mean it would have to be 0? is that what you're getting at?

OpenStudy (zarkon):

subtract \[\begin{array}{c} 2a+b=0\\ 3a+b=0\\\hline -a=0 \end{array}\] thus a=0 if a=0 then \(2a+b=0\Rightarrow 2(0)+b=0\Rightarrow b=0\)

OpenStudy (anonymous):

oh okay so that shows that they must be independent. Thanks!